# Human-in-the-Loop Local Corrections of 3D Scene Layouts via Infilling

### v1

@@ Abstract
Figure 1: Starting with an initial layout (top left) predicted by our model in global prediction mode, our system allows a user to identify local errors within the layout. The user then prompts our model in local correction mode to refine these entities, resulting in improved accuracy. The user repeats this process (top center) until they achieve the desired layout (top right). (bottom) To perform the local correction, our method masks out the selected entities and infills them with the refined entities. We highlight an incorrect wall structure in the initial layout (pink circle) that is substantially improved (green circle) in the refined layout.
@@ 1 Introduction
While interactive methods have seen success in 2D computer vision, such as with the Segment Anything Model (SAM) [21], applying a human-in-the-loop paradigm to 3D tasks remains largely unexplored. To explore human-in-the-loop paradigms in 3D, we focus on 3D scene layout estimation conducted from an egocentric context. This task provides an effective starting point for studying how human input can prompt model-driven local refinements to improve 3D predictions. We build upon SceneScript [3], a promising technique for scene layout estimation that leverages language modeling to perform a single-shot global prediction of 3D scene layout from an egocentric walkthrough.
To perform this iterative refinement for 3D scene layout, we introduce the novel task of local correction where a learned model must fix a local error in the scene layout estimate, which is identified by a user. We adapt SceneScript to simultaneously learn the task of local correction in addition to global prediction. By framing this task as ”infilling” [13, 15, 29, 4], which aims to generate text to fill in missing spans of a document, we leverage insights from natural language processing to develop a multi-task SceneScript model that can perform both global predictions and local corrections. Interestingly, this multi-task approach maintains strong performance on global predictions compared to the original SceneScript model, while yielding significantly stronger results on local corrections.
We integrate our model into a human-in-the-loop system that allows a user to seamlessly refine scene layout estimates via a low-friction “one-click fix” workflow. Our system initializes the scene layout estimate using the model’s global prediction. Then, the user identifies (“one-click”) an error and triggers the same model’s local correction capability (“fix”) to refine it. This process is iterated until the desired layout is achieved. To immersively engage with 3D content, we utilize mixed reality technologies via a Meta Quest 3 headset with a Project Aria [31] rig rigidly attached on top. Figure 1 visualizes our interactive system. Our iterative local corrections allow the final refined layout to diverge from the training distribution, which is crucial in modeling complex real-world scenes.
@@
  • We integrate our model into a human-in-the-loop system in a mixed reality environment employing a “one-click fix” workflow for seamless iterative refinements, enabling the final refined layout to diverge from the training distribution and effectively model complex real-world scenes.
@@ 3 Method
We use SceneScript’s [3] structured language commands to represent scene layout. A hand-designed set of commands and parameters is used to describe basic architectural entities for scene layout, e.g. make_door with parameters such as height, width. A scene layout can be described as a sequence of these commands. This can then be converted into a sequence of tokens which is ready to be processed by a language model. For this work, we only utilize architectural geometries, i.e. walls, doors, and windows, and do not use bounding boxes nor volumetric primitives. Extension to those additional commands is straightforward.
We employ SceneScript’s encoder-decoder architecture. The model consumes visual data as input and outputs a sequence of tokens that can be de-serialized into structured language commands. We use the version of SceneScript that encodes the semi-dense point clouds only. The encoder consists of a sparse 3D ResNet [33] that produces features which are cross-attended to by a decoder-only Transformer [36]. Note that we do not introduce any changes to this architecture. Please refer to [3] for more details about the structured language commands, custom tokenization procedure, and model details.
@@ 4.2 Datasets, Metrics, and Baselines
Due to the novelty of both task and system, there are no existing baselines for either. For local correction, we created a baseline using vanilla SceneScript (i.e., trained only on global predictions) via partial layout completion [18]: given all entities except the ones to correct, continue generating tokens until <STOP>. However, to our knowledge, no human-in-the-loop systems exist for low-friction editing of 3D scene layouts.
@@ 4.4.1 Automatic Heuristic Local Corrections
Starting from the same initial layout estimate, we run the loop for 5 trials with 50 iterations each, and show Average PlaneDistance (lower is better) results averaged over the 5 trials, all 25 AEO sequences, and all classes (wall/door/window). Because the real-world problem is much harder than simulation, we opt for PlaneDistance which is simply the distance function $d_{E}(\cdot)$ defined in [3], as opposed to AvgF1. At every iteration, we compute PlaneDistance of the current layout compared to ground truth. In Figure 3, we show results of a SceneScript baseline and our model trained with different language orderings integrated the iterative system. All of our multi-task models show a trend of improvement as opposed to the baseline, which stays relatively flat.
@@ 4.4.2 Human-in-the-Loop System
As there is no existing baseline system, we conducted a user study comparing the integration of the vanilla SceneScript baseline vs. our model in our human-in-the-loop system. Please refer to the supplement for details and results.
@@ 5 Discussion and Limitations
Our approach demonstrates strong performance in predicting and correcting room structures, such as walls, doors, and windows. Because we build on top of SceneScript’s structured language representation, we can easily adapt to new spatial representations, including 3D bounding boxes and volumetric primitives, by updating the language schema. By design, we do not use selected entities as input during local correction, thus we do not need special augmentations for each class. Furthermore, we have only tackled a single new task framed as infilling. It would be valuable to explore how to frame additional tasks as infilling, such as updating object movements based on the latest observations to an existing map, and integrate them into our framework.
@@ A.2 Multi-Wall Evaluation
In Tables 9 and 10, we show an analysis of this breakdown by $n_{w}$ for our models trained with different language orderings on ASE [3] and our internal version of ASE, respectively. We separate the evaluation from 1 to $\geq 4$ walls. In both tables, performance on $n_{w}=1$ is nearly perfect as the problem is easy. However, the performance of all models degrades as $n_{w}$ increases, because the problem becomes more difficult. Similarly, the performance gaps for the methods generally increase as the number of walls increases.

### v2

@@ Abstract
Figure 1: Starting with an initial layout (top left) predicted by our model in global prediction mode, our system allows a user to identify local errors within the layout. The user then prompts our model in local correction mode to refine these entities, resulting in improved accuracy. The user repeats this process (top center) until they achieve the desired layout (top right). (bottom) To perform the local correction, our method masks out the selected entities and infills them with the refined entities. We highlight an incorrect wall structure in the initial layout (pink circle) that is substantially improved (green circle) in the refined layout. Please see our project webpage for videos of our system.
@@ 1 Introduction
While interactive methods have seen success in 2D computer vision, such as with the Segment Anything Model (SAM) [21], applying a human-in-the-loop paradigm to 3D tasks remains largely unexplored. To explore human-in-the-loop paradigms in 3D, we focus on 3D scene layout estimation conducted from an egocentric context. We build upon SceneScript [3], a promising technique for scene layout estimation that leverages language modeling to perform a single-shot global prediction of 3D scene layout from an egocentric walkthrough.
To perform this iterative refinement for 3D scene layout, we introduce the novel task of local correction where a learned model must fix a local error in the scene layout estimate, which is identified by a user. We adapt SceneScript to simultaneously learn the task of local correction in addition to global prediction. By framing this task as ”infilling” [13, 15, 29, 4], which aims to generate text to fill in missing spans of a document, we leverage insights from natural language processing to develop a multi-task SceneScript model that can perform both global predictions and local corrections. Interestingly, this multi-task approach maintains strong performance on global predictions compared to the original SceneScript model, while yielding significantly stronger results on local corrections. We focus this investigation on architectural entities (walls/doors/windows) to prove the core concept of “infilling”. As our method leverages SceneScript’s [3] structured language, it can be easily extended to more general 3D scene layout including bounding boxes and/or volumetric primitives.
We integrate our model into a human-in-the-loop system that allows a user to seamlessly refine scene layout estimates via a low-friction “one-click fix” workflow. Our system initializes the scene layout estimate using the model’s global prediction. Then, the user identifies (“one-click”) an error and triggers the same model’s local correction capability (“fix”) to refine it. This process is iterated until the desired layout is achieved. To immersively engage with 3D content, we utilize mixed reality technologies via a Meta Quest 3 headset with a Project Aria [31] rig rigidly attached on top. Figure 1 visualizes our interactive system (see supplemental video). Our iterative local corrections allow the final refined layout to diverge from the training distribution, which is crucial in modeling complex real-world scenes.
@@
  • We integrate our model into a human-in-the-loop system in a mixed reality environment employing a “one-click fix” workflow for seamless iterative refinements, enabling the final refined layout to diverge from the training distribution and effectively model complex real-world scenes. Please see our project webpage for videos of our system.
@@ 3 Method
We use SceneScript’s [3] structured language commands to represent scene layout. A hand-designed set of commands and parameters is used to describe geometric entities for scene layout, e.g. make_door with parameters such as height, width. A scene layout can be described as a sequence of these commands, which can be converted into a sequence of tokens ready to be processed by a language model. To investigate the core concept of “infilling”, we focus only on architectural geometries in this work, i.e. walls, doors, and windows. Our method is agnostic to any structured language definition, thus it can directly extended to other commands (e.g. bounding boxes) via plug-and-play.
We employ SceneScript’s encoder-decoder architecture. The model consumes semi-dense point clouds (no images) and outputs a sequence of tokens that can be de-serialized into structured language commands. The encoder is a sparse 3D ResNet [33] that produces features which are cross-attended to by a Transformer Decoder [36]. Note that we do not introduce any changes to this architecture. Please refer to [3] for details about the structured language commands, custom tokenization procedure, and model details.
@@ 4.2 Datasets, Metrics, and Baselines
Due to the novelty of both task and system, there are no existing baselines for either. For local correction, we created a baseline using vanilla SceneScript (trained only on global predictions) via partial layout completion [18]: given all entities except the ones to correct, continue generating tokens until <STOP>. To our knowledge, no human-in-the-loop systems exist for low-friction editing of scene layouts.
@@ 4.4.1 Automatic Heuristic Local Corrections
Starting from the same initial layout estimate, we run the loop for 5 trials with 50 iterations each, and show Average PlaneDistance (lower is better) results averaged over the 5 trials, all 25 AEO sequences, and all classes (wall/door/window). Because the real-world problem is harder than simulation, we opt for PlaneDistance, i.e. the distance function $d_{E}(\cdot)$ defined in [3], as opposed to AvgF1. At every iteration, we compute PlaneDistance of the current layout compared to ground truth. In Figure 3, we show results of a SceneScript baseline and our model trained with different language orderings integrated the iterative system. All of our multi-task models show a trend of improvement as opposed to the baseline, which stays relatively flat.
@@ 4.4.2 Human-in-the-Loop System
As there is no existing baseline system, we conducted a user study comparing the integration of the vanilla SceneScript baseline vs. our model in our human-in-the-loop system. Please refer to the supplement for details and results, and the project webpage for videos of our system.
@@ 5 Discussion and Limitations
Our approach demonstrates strong performance in predicting and correcting room structures. Because we build on top of SceneScript’s structured language, we can easily adapt to new structures, including 3D bounding boxes and volumetric primitives, by updating the language schema. By design, we do not use selected entities as input during local correction, facilitating simple plug-and-play extensions.
Existing semi-automatic layout annotation tools, such as PanoAnnotator [39], rely on continuous-value edits that demand significant user effort. Our method instead introduces a lower-friction “one-click fix” paradigm. Enabling users to correct geometric errors with a single click is non-trivial, posing an interesting problem our work seeks to address.
@@ A.2 Multi-Wall Evaluation
In Tables 9 and 10, we show an analysis of this breakdown by $n_{w}$ for our models trained with different language orderings on ASE [3] and our internal version of ASE, respectively. We separate the evaluation from 1 to 4$\geq$ 4≥ 4 walls. In both tables, performance on $n_{w}=1$ is nearly perfect as the problem is easy. However, the performance of all models degrades as $n_{w}$ increases, because the problem becomes more difficult. Similarly, the performance gaps for the methods generally increase as the number of walls increases.
@@ E.1 Global Prediction with Accumulated Points
E.2 Demo Videos
At our project webpage, we provide videos of our human-in-the-loop system running live on a Meta Quest 3 with an Aria device rigidly attached. The three videos correspond to the three real-world scenes used for qualitative results in the main paper (one from Figure 1 and two from Figure 4). On the left of the video, we highlight the current layout estimate, action, selected entities, and infilled entities.
E.3 Failure Cases
In our supplemental materials, we provide a video of failure cases of our human-in-the-loop system. We detail them here:
  • Our model struggles to accurately identify individual windows when multiple small windows are positioned closely together. Groups of windows are often misidentified as a single large window. We posit that this is due to the nature of the synthetic training data. This can also be seen in our qualitative results in Figure 4.
  • Although users typically expect a newly added entity to appear directly in front, our model does not consistently follow this behavior in practice. This is likely due to the heuristics used to select data (e.g. visibility as discussed in Section 3.1.2). Future work involves curating the data to better align the model’s behavior with human expectations.
  • Lastly, we show an example where the model is not capable of producing the correct geometry for a pair of small windows. We speculate that this style of window is not present anywhere in our synthetic training data. Although the local correction task demonstrates better generalization when used in a human-in-the-loop system compared to the global prediction distribution, it is inherently limited to the learned local distributions. As a result, it cannot predict outcomes for patterns or scenarios that fall outside these distributions.